A solution is made by dissolving 513 grams of sucrose (C12H22O11) in 2 liters of water. Calculate the mole fraction of each component in the solution. Molar mass of sucrose = 342 g/mol, water = 18 g/mol.
1
Expert's answer
2020-09-29T06:46:54-0400

 513 grams of sucrose Have moles = 513 / 342 = 1.5 moles .


 2 liters of water have 2000 ml of wate .


density of water is 1 g / ml .


so , mass of water = 2000 g .


moles of water = 2000 / 18 = 111.11 mol .


So , mole fraction of sucrose = moles of sucrose / total moes of solution .

= 1.5 / 1.5 + 111.11


= 0.013320 .



mole fraction of water = 1 - mole fraction of sucrose = 1- 0.013320 = 0.98667